OpenStudy (anonymous):

If 3x = 0, why cant we take x to the other side, then, 3=0, which is not true? WHats wrong in this????????

OpenStudy (anonymous):

If three x is zero.. then that means what ever X is is zero because something times zero is always zero.

hartnn (hartnn):

how are you planning to take x on the other side ? multiplying, dividing ?

OpenStudy (anonymous):

3 is still three :)

OpenStudy (anonymous):

If you were to take it to the other side.. it would have to be divided.

OpenStudy (anonymous):

by dividing

hartnn (hartnn):

you know you cannot divide by a 0 ? thats illegal...

hartnn (hartnn):

3x = 0 since 3 is not a 0, x=0 and so, you cannot divide by 0.

OpenStudy (anonymous):

i am not dividing by 0. i m dividing by x

hartnn (hartnn):

but x here is actualy =0

OpenStudy (anonymous):

yeah... i noe but there is nothing wrong in these steps

OpenStudy (anonymous):

but the result is wrong!

ganeshie8 (ganeshie8):

u cant divide by x, without having full knowledge of wat values it can take

OpenStudy (anonymous):

Haha that's impossible if you think about it mathematically.. if i have three 'nothings' and i were to give 3 people 1 each- everyone would have nothing...

ganeshie8 (ganeshie8):

remember x is a variable, it can take any value

OpenStudy (anonymous):

for 3x = 5xy, we dont know the value of x but still we can divide by x and write it as 3 = 5y or 7x = 5y can be written as 7 = 5y/x

hartnn (hartnn):

not actually, even thats incomplete

ganeshie8 (ganeshie8):

whenever u divide by x, put an explicit caveat saying x \(\ne\) 0

hartnn (hartnn):

3x = 5xy 3x-5xy =0 x (3-5y) = 0 x=0 or 3-5y =0 this is correct way

hartnn (hartnn):

so, when 3-5y =0 then x is NOT = 0!!

hartnn (hartnn):

notice the "or" in between either of it can be true

hartnn (hartnn):

7x = 5y y = 5y/x <----incomplete y=5y/x \(x \ne 0\)<---complete

OpenStudy (anonymous):

OMG its so complicated

hartnn (hartnn):

but logical, since we cannot divide by 0, we need to mention :)

OpenStudy (anonymous):

let it just be like 3 is a constant therefore it cant be equal to 0

hartnn (hartnn):

thats what we said.

OpenStudy (anonymous):

yeah... :)